# Distributional Evaluation of Generative Models via Relative Density Ratio

### v1

@@ Abstract
We propose a functional evaluation metric for generative models based on the relative density ratio (RDR) designed to characterize distributional differences between real and generated samples. We show that the RDR as a functional summary of the goodness-of-fit for the generative model, possesses several desirable theoretical properties. It preserves $\phi$-divergence between two distributions, enables sample-level evaluation that facilitates downstream investigations of feature-specific distributional differences, and has a bounded range that affords clear interpretability and numerical stability. Functional estimation of the RDR is achieved efficiently through convex optimization on the variational form of $\phi$-divergence. We provide theoretical convergence rate guarantees for general estimators based on M-estimator theory, as well as the convergence rates of neural network-based estimators when the true ratio is in the anisotropic Besov space. We demonstrate the power of the proposed RDR-based evaluation through numerical experiments on MNIST, CelebA64, and the American Gut project microbiome data. We show that the estimated RDR not only allows for an effective comparison of the overall performance of competing generative models, but it can also offer a convenient means of revealing the nature of the underlying goodness-of-fit. This enables one to assess support overlap, coverage, and fidelity while pinpointing regions of the sample space where generators concentrate and revealing the features that drive the most salient distributional differences.
@@ 1 Introduction
Generative AI has become an integral part of modern society, facilitating tasks in everyday life and advancing scientific discovery. With the remarkable development and deployment of generative models in domains ranging from text and image generation to scientific data synthesis, there is growing interest in determining which models perform better in different contexts. We have witnessed great success in text and image generation using various model classes, including VAEs, GANs, diffusion models, and continuous flow matching models. However, despite the realism of generated samples to the human eye, studies have revealed distributional shifts between samples produced by advanced generators and those observed in real data. This phenomenon has been observed for many modern generators, such as GANs (Richardson and Weiss, 2018) and diffusion models (Sehwag et al., 2022), which are not easily detected by visual inspection. Such discrepancies are particularly concerning when generative AI is used for scientific applications, where the fidelity of the underlying data distribution is critical. Thus, a key step toward trustworthy generative modeling is to understand, at a distributional level, how and where the generator diverges from the true data generation mechanism.
@@
Existing automated evaluation criteria for modern generative models have largely focused on benchmarking generators through scalar metrics. Popular metrics include the Fréchet Inception Distance (Heusel et al., 2017, FID) and the Inception Score (Salimans et al., 2016, IS), both of which rely on strong parametric assumptions or low-dimensional feature mappings to assess model performance on large-scale, high-dimensional data. Simon et al. (2019); Sajjadi et al. (2018) provides a precision-recall curve to describe the coverage and fidelity of generators based on low-dimensional features of the original data. While these metrics capture human-perceivable fidelity, they largely overlook underlying distributional discrepancies. As a result, even a generator that samples extensively from a narrow region of the true data support can achieve a high score, provided that the generated samples look realistic. More importantly, in scientific data generation, as we will demonstrate in a microbiome example, visually convincing generations may be biased and fail to represent the true data-generating mechanism in important ways that could bias the downstream scientific analysis. This issue is critical in many scientific application domains where the generated samples are used to augment observed data (Choi et al., 2023) and to perform downstream analyzes such as biomarker detection (Marouf et al., 2020) or cell clustering (Lall et al., 2022). The goal of this work is to propose a functional metric, rather than a single scalar metric, capable of providing both distribution-level and sample-level assessments. Our approach is based on the relative density ratio, which is simple to implement and scalable to massive, high-dimensional datasets. We illustrate how this framework allows for the identification of features associated with the detected distributional discrepancies between observed and generated data.
@@
The classic two-sample comparison methods in statistics have mostly adopted a hypothesis testing perspective. The predominant approach involves computing a test statistic, which is usually a scalar distance metric, and based on that, computing the p-value under its (asymptotic) null distribution. Examples of univariate samples include the Kolmogorov–Smirnov test and the Anderson-Darling test (Razali et al., 2011). Modern nonparametric methods, such as Maximum Mean Discrepancy (MMD) and Energy Distance (Gretton et al., 2012; Székely and Rizzo, 2013), use prespecified kernel embeddings or Euclidean distances to provide a distance metric between two distributions. These methods aim to address the question of how far two distributions are apart through a single scalar metric. As such, they cannot reveal exactly how and where they differ. One is often satisfied with a simulation based $p$-value for these metrics attained through resampling.
@@
In general, we question the relevance of the hypothesis testing perspective on two-sample comparisons in the context of assessing modern generative models—all models are wrong, and all generative models are definitely quite wrong in applications that involve highly complex data generative mechanisms. There is no question that the null hypothesis, with sufficient data (often available in synthetic experiments), will be rejected. The more valuable question is how the generative model fails to capture the truth. As such, we believe the functional estimation perspective is more appropriate.
@@
Which function should we estimate as a way to summarize two-sample discrepancies? A natural candidate is the density ratio, i.e., the ratio of the two sampling densities. Indeed, density ratio estimation (DRE) has been studied, and well-known approaches include kernel-based methods such as KLIEP (Sugiyama et al., 2007) and uLSIF (Kanamori et al., 2009). More recently, Awaya et al. (2025) introduced an approach that uses additive tree ensembles along with a new loss function to learn the DR and achieves generalized Bayesian uncertainty quantification. A fundamental assumption of the DRE approach is that the two distributions must have the same support; otherwise, the ratio is undefined in parts of the space. Such an assumption is easily violated in very high-dimensional, complex data generative settings. This issue can incur severe extrapolation errors, a form of overfitting, when highly flexible function approximators, such as neural-networks, are adopted for DRE.
@@
as our choice of the functional metric, which Yamada et al. (2013) first considered and referred to as the relative density ratio (RDR). In our setting, the denominator corresponds specifically to the 50–50 mixture of $p$ and $q$. While Yamada et al. (2013) is largely grounded in kernel-based methods for RDR estimation, our estimation approach is tailored to neural-network-based estimators using $\phi$-divergence objectives, following the general strategy introduced in Nguyen et al. (2010). We will show in Section 2 that the RDR has a bounded range, good interpretability, and numerical stability for neural network estimators.
Alongside the proposed relative density ratio framework, we contribute new theoretical insights and empirical results on both image generation and scientific domains, including the relative abundance microbiome data from the American Gut Project (McDonald et al., 2018b). We first show that the ratio estimator can be viewed as an M-estimator, and we establish consistency and convergence rate results for general estimator classes that satisfy certain entropy bounds. Specifically, for the sparse neural network estimator, building on prior work by Suzuki and Nitanda (2021), we show that if the true ratio function lies in an anisotropic Besov space, a function space that allows differing degrees of smoothness across coordinates in high-dimensional settings, then the convergence rate of the neural network estimator is nearly optimal up to a logarithmic factor in $n$, and this rate is dimension-free.
In our numerical experiments, we compare a range of modern generative models, including variational autoencoders (VAEs), generative adversarial networks (deep convolutional GANs, Radford et al., 2015), diffusion models (DDIM, Song et al., 2021b), and continuous flow matching models (ICFM, Lipman et al., 2022), on image datasets such as MNIST and CelebA-64, as well as on the American Gut data. We found that even for the well-performing generators with high fidelity and good coverage, in a simple task of generating digits from MNIST data, the generator tends to have its own bias and concentrate on certain local regions in the high-dimensional support. A similar trend also appears in the higher-dimensional CelebA-64 data and the more complex, highly sparse microbiome data. Furthermore, in the experiments of CelebA-64 and the American Gut Project, we demonstrate how to use this ratio function to detect the features or attributes associated with such distributional discrepancies, making it more interpretable.
@@ 2 Distributional Evaluation via Relative Density Ratio
The RDR $r(x)$ has a bounded image $[0,2]$. The squared Hellinger distance between $p(x)$ and $\frac{1}{2}\left\{p(x)+q(x)\right\}$ is bounded above by $1-\frac{1}{\sqrt{2}}(\approx 0.293)$.
@@
Note that $r=T(g)$, where $T(g)=\frac{2g}{g+1}$, which is a one-to-one mapping between $r(x)$ and $g(x)$. Because of the one-to-one mapping, Theorem 1 indicates that we can use the 1D empirical distributions formed by $\left\{r(X_{i})\right\}_{i=1}^{n},\left\{r(Y_{j})\right\}_{j=1}^{m}$ to summarize the distributional difference between $P$ and $Q$ in high-dimensions while preserving the $\phi$-divergence $D_{\phi}(P\|Q)$. We can use the 1D histograms of $\left\{r(X_{i})\right\}_{i=1}^{n},\left\{r(Y_{j})\right\}_{j=1}^{m}$ as the distributional visualization of the original two samples; histograms such as Row 3 in Figure 2 demonstrate the coverage/fidelity of the generator of high-dimensional samples.
We first introduce the estimation method of the ratio function $r(x)$ in Subsection 2.1, and then illustrate the usage of RDR in Subsection 2.3 with simple 1D examples in connection with existing two-sample comparison methods.
@@ 2.1 Estimation via the variational form of $\phi$-divergence
Our estimation of $r(x)$ is closely related to the density ratio $g(x)=p(x)/q(x)$. If $P\ll Q$ and $\phi:\mathbb{R}^{+}\to\mathbb{R}$ is a convex function with $\phi(1)$=0, let $D_{\phi}$ be the general $\phi$-divergence, defined as $D_{\phi}(P\|Q):=\int_{\mathcal{X}}\phi\left(\frac{\text{d}P}{\text{d}Q}(x)\right)\text{d}Q(x)$. Table 1 lists a few common options for $\phi$. The variational form (dual form) of $\phi$-divergence provides a way to learn the density ratio $g(x)$. This theoretical property of $\phi$-divergence is proved in Nguyen et al. (2010). In convex optimization (Rockafellar, 1997), the subdifferential of a convex function $\phi$ at point $t\in\mathbb{R}$ is defined as the set
@@
We reiterate Lemma 1 in Nguyen et al. (2010) since it is the foundation of our estimation procedure.
Lemma 1 (Nguyen et al. (2010)).
For any $\phi$-divergence, the dual form can be written as an optimization problem with respect to a functional class $\mathcal{F}$ mapping from $\mathcal{X}$ to $\mathbb{R}$,
@@
This result indicates that the optimizer of the variational form of $\phi$ divergences is a certain function of the density ratio $g$, leading to many ways of estimating the density ratio. Table 1 shows the optimizers for KL-divergence, chi-squared divergence, and the squared Hellinger divergence. In fact, this idea has been explored by Nowozin et al. (2016) to use the density ratio as the discriminator in a GAN (Generative Adversarial Network).
@@ 2.2 A neural network model to estimate the (relative) density ratio
When $X_{1},\dots,X_{n}\overset{\scriptsize\mbox{iid}}{\sim}P$, and $Y_{1},\dots,Y_{m}\overset{\scriptsize\mbox{iid}}{\sim}Q$, for any measurable set $A\subset\mathcal{X}$, denote the empirical distributions $P_{n}(A):=\frac{1}{n}\sum_{i=1}^{n}\bm{1}\left\{X_{i}\in A\right\}$, and $Q_{m}(A):=\frac{1}{m}\sum_{j=1}^{m}\bm{1}\left\{Y_{j}\in A\right\}$. Using a change-of-variable transformation $g=(1-2f)^{-2}$ from the variational form of the squared Hellinger distance in Table 1, let $\hat{g}_{n}$ be the estimated maximizer over a functional class of candidates $\mathcal{G}$, based on the observed samples,
@@
The estimation method depends on the class of candidate functions in $\mathcal{G}$ and the smoothness of the true function $g_{0}$. Awaya et al. (2025) considers additive tree ensembles for $\log g$ as their $\mathcal{G}$. Note that $g^{*}=p/q$ is the unique global minimizer of $\frac{1}{2}\int g^{-\frac{1}{2}}\text{d}P+\frac{1}{2}\int g^{\frac{1}{2}}\text{d}Q$. In this work, we focus on using neural networks to estimate the ratio functions under this loss through gradient-based optimization. In practice, (stochastic) gradient descent algorithms to find the optimal $\hat{g}_{n},\hat{r}_{n}$.
The feedforward neural network can be viewed as a functional learner (Sazlı, 2006), parameterized with weight matrices $\left\{W^{(l)}\right\}_{l=1}^{L}$ and bias vector $\left\{b^{(l)}\right\}_{l=1}^{L}$. Let $U_{l}$ be the total number of nodes in layer $l$ when $1<l\leq L$, $W^{(l)}\in\mathbb{R}^{U_{l}\times U_{l-1}}$, and the vector-valued activation function $\sigma^{(l)}:\mathbb{R}^{U_{l}}\to\mathbb{R}^{U_{l}}$ is an element-wise Lipschitz function. Denote $\bm{\theta}:=\left\{W^{(l)},b^{(l)}\right\}_{l=1}^{L}$ as the set of all parameters, a neural network estimator can be formulated as a sequence of compositions
@@
The empirical loss being minimized was first introduced in Awaya et al. (2025) and referred to as the “balancing loss” for reasons explained in that paper. The optimizer in (3) can be estimated by $\hat{g}_{n}(x)=f(x;\widehat{\bm{\theta}}_{g})$. Similarly, given the objective in (4) and the learned $\widehat{\bm{\theta}}_{r}$, $\hat{r}_{n}=f(x;\widehat{\bm{\theta}}_{r})$. The neural network architecture varies in different data applications: for image data, variations of convolutional neural networks (CNNs) are applied, and for lower-dimensional examples and microbiome data, multilayer perceptrons (MLPs) are sufficient. Later in Section 3, we will first provide a general convergence theory for general estimators with constraints on $\mathcal{G}$, followed by a convergence rate result using the sieved sparse neural networks space as the functional class $\mathcal{G}_{n}$, where the size of the neural network space is allowed to grow with the sample size $n$.

### v2

@@ Abstract
We propose a function-valued evaluation metric for generative models based on the relative density ratio (RDR) designed to characterize distributional differences between real and generated samples. As an evaluation metric, the RDR function preserves $\phi$-divergence between two distributions, enables sample-level evaluation that facilitates downstream investigations of feature-specific distributional differences, and has a bounded range that affords clear interpretability and numerical stability. Function estimation of the RDR is achieved efficiently through optimization on the variational form of $\phi$-divergence. We provide theoretical convergence rate guarantees for general estimators based on M-estimator theory, as well as the convergence rate of neural network-based estimators when the true ratio is in the anisotropic Besov space. We demonstrate the power of the proposed RDR-based evaluation through numerical experiments on MNIST, CelebA64, and the American Gut project microbiome data. We show that the estimated RDR enables not only effective overall comparison of competing generative models, but also a convenient way to reveal the underlying nature of goodness-of-fit. This enables one to assess support overlap, coverage, and fidelity while pinpointing regions of the sample space where generators concentrate and revealing the features that drive the most salient distributional differences.
@@ 1 Introduction
Generative AI has become an integral part of modern society, facilitating tasks in everyday life and advancing scientific discovery. With the remarkable development and deployment of generative models in domains ranging from text and image generation to scientific data synthesis, there is a growing interest in determining which models perform better in different contexts. We have witnessed great successes in text and image generation using various model classes, including VAE (variational autoencoders), GANs (generative adversarial networks), diffusion models, and continuous flow matching models. However, despite the realism of generated samples to the human eye, studies have revealed distributional shifts between samples produced by advanced generators and those observed in real data. This phenomenon has been observed for many modern generators, such as GANs (richardson2018gans) and diffusion models (sehwag2022generating), which are not easily detected by visual inspection. Such discrepancies are particularly concerning when generative AI is used for scientific applications, where the fidelity of the underlying data distribution is critical. Thus, a key step toward trustworthy generative modeling is to understand, at a distributional level, how and where the generator diverges from the true data generation mechanism.
@@
Existing automated evaluation criteria for modern generative models have largely focused on benchmarking generators through scalar metrics. Popular metrics include the Fréchet Inception Distance (heusel2017gans, FID) and the Inception Score (salimans2016improved, IS), both of which rely on strong parametric assumptions or low-dimensional feature mappings to assess model performance on large-scale, high-dimensional data. pmlr-v97-simon19a and sajjadi2018assessing provide a precision-recall curve to describe the coverage and fidelity of generators based on low-dimensional features of the original data. While these metrics capture human-perceivable fidelity, they largely overlook underlying distributional discrepancies. As a result, even a generator that samples extensively from a narrow region of the true data support can achieve a high score, provided that the generated samples look realistic. More importantly, in scientific data generation, as we will demonstrate in a microbiome example, visually convincing generations may be biased and fail to represent the true data-generating mechanism in important ways that could bias the downstream scientific analysis. This issue is critical in many scientific application domains where the generated samples are used to augment observed data (choi2023deepmicrogen) and to perform downstream analyzes such as biomarker detection (marouf2020realistic) or cell clustering (lall2022lsh). The goal of this work is to propose a function-valued metric, rather than a single scalar metric, capable of providing both distribution-level and sample-level assessments. Our approach is based on the relative density ratio, which is simple to implement and scalable to massive, high-dimensional datasets. We illustrate how this framework allows for the identification of features associated with the detected distributional discrepancies between observed and generated data.
@@
Comparing the generative model distribution and the real data distribution can be viewed as the classic two-sample problem from a statistical perspective. The classic two-sample comparison methods in statistics have mostly adopted a hypothesis testing perspective. The predominant approach involves computing a test statistic, which is usually a scalar distance metric, and based on that, computing the p-value under its (asymptotic) null distribution. Examples of univariate samples include the Kolmogorov–Smirnov test and the Anderson-Darling test (razali2011power). Modern nonparametric methods, such as Maximum Mean Discrepancy (MMD) and Energy Distance (gretton2012kernel; szekely2013energy), use prespecified kernel embeddings or Euclidean distances to provide a distance metric between two distributions. These methods aim to address the question of how far two distributions are apart through a single scalar metric. As such, they cannot reveal exactly how and where they differ. One is often satisfied with a simulation based $p$-value for these metrics attained through resampling.
@@
In general, we question the relevance of the hypothesis testing perspective on two-sample comparisons in the context of assessing modern generative models—all models are wrong, and all generative models are definitely quite wrong in applications that involve highly complex data generative mechanisms. There is no question that the null hypothesis, with sufficient data (often available in synthetic experiments), will be rejected. The more valuable question is how the generative model fails to capture the truth. As such, we believe the function estimation perspective is more appropriate.
@@
Which function should we estimate as a way to summarize two-sample discrepancies? A natural candidate is the density ratio, i.e., the ratio of the two sampling densities. Indeed, density ratio estimation (DRE) has been studied, and well-known approaches include kernel-based methods such as KLIEP (sugiyama2007direct) and uLSIF (kanamori2009least). More recently, awaya2025two introduced an approach that uses additive tree ensembles along with a new loss function to learn the DR and achieves generalized Bayesian uncertainty quantification. A fundamental assumption of the DRE approach is that the two distributions must have the same support; otherwise, the ratio is undefined in parts of the space. Such an assumption is easily violated in very high-dimensional, complex data generative settings. This issue can incur severe extrapolation errors, a form of overfitting, when highly flexible function approximators, such as neural-networks, are adopted for DRE.
@@
as our choice of the function-valued metric, which yamada2013relative first considered and referred to as the relative density ratio (RDR). In our setting, the denominator corresponds specifically to the 50–50 mixture of $p$ and $q$. While yamada2013relative is largely grounded in kernel-based methods for RDR estimation, our estimation approach is tailored to neural-network-based estimators using $\phi$-divergence objectives, following the general strategy introduced in Nguyen2010-xb. We will show in Section 2 that the RDR has a bounded range, good interpretability, and numerical stability for neural network estimators.
Alongside the proposed relative density ratio framework, we contribute new theoretical insights and empirical results on both image generation and scientific data domains, including the relative abundance microbiome data from the American Gut Project (AmericanGut). We first show that the ratio estimator can be viewed as an M-estimator, and we establish consistency and convergence rate results for general estimator classes that satisfy certain entropy bounds. Specifically, for the sparse neural network estimator, building on prior work by suzuki2021deep, we show that if the true ratio function lies in an anisotropic Besov space, a function space that allows differing degrees of smoothness across coordinates, then the convergence rate of the neural network estimator is nearly optimal up to a logarithmic factor in $n$, and depends only on the smoothness of the true ratio function, rather than directly on the sample dimension.
In our numerical experiments, we compare a range of modern generative models, including variational autoencoders (VAEs), generative adversarial networks (deep convolutional GANs, DCGAN), diffusion models (DDIM, song2021ddim), and continuous flow matching models (ICFM, lipman2022flow), on image datasets such as MNIST and CelebA-64, as well as on the American Gut data. We found that even for the well-performing generators with high fidelity and good coverage, in a simple task of generating digits from MNIST data, the generator tends to have its own bias and concentrate on certain local regions in the high-dimensional support. A similar trend also appears in the higher-dimensional CelebA-64 data and the more complex, highly sparse microbiome data. Furthermore, in the experiments of CelebA-64 and the American Gut Project, we demonstrate how to use this ratio function to detect the features or attributes associated with such distributional discrepancies, making it more interpretable.
@@ 2 Distributional Evaluation via Relative Density Ratio
The RDR $r(x)$ has a bounded image $[0,2]$. The squared Hellinger distance between the numerator density $p(x)$ and the denominator density $\frac{1}{2}\left\{p(x)+q(x)\right\}$ is bounded above by $1-\frac{1}{\sqrt{2}}(\approx 0.293)$.
@@
Note that there is a one-to-one mapping between RDR and the density ratio: $r=T(g)$, where $T(g)=\frac{2g}{g+1}$, which is a one-to-one mapping between $r(x)$ and $g(x)$. Because of the one-to-one mapping, Theorem 1 indicates that we can use the 1D empirical distributions formed by $\left\{r(X_{i})\right\}_{i=1}^{n},\left\{r(Y_{j})\right\}_{j=1}^{m}$ to summarize the distributional difference between $P$ and $Q$ in high-dimensions while preserving the $\phi$-divergence $D_{\phi}(P\|Q)$. We can use the 1D histograms of $\left\{r(X_{i})\right\}_{i=1}^{n}$, $\left\{r(Y_{j})\right\}_{j=1}^{m}$ as the distributional visualization of the original two samples; histograms such as Row 3 in Figure 2 demonstrate the coverage/fidelity of the generator of high-dimensional samples.
We first introduce the estimation method of the ratio function $r(x)$ in Section 2.1, and then illustrate the usage of RDR in Section 2.3 with simple 1D examples in connection with existing two-sample comparison methods.
@@ 2.1 Estimation via the variational form of $\phi$-divergence
Our estimation of $r(x)$ is closely related to the density ratio $g(x)=p(x)/q(x)$. If $P\ll Q$ and $\phi:\mathbb{R}^{+}\to\mathbb{R}$ is a convex function with $\phi(1)$=0, let $D_{\phi}$ be the general $\phi$-divergence, defined as $D_{\phi}(P\|Q):=\int_{\mathcal{X}}\phi\left(\frac{\text{d}P}{\text{d}Q}(x)\right)\text{d}Q(x)$. Table 1 lists a few common options for $\phi$. The variational form (dual form) of $\phi$-divergence provides a way to learn the density ratio $g(x)$. This theoretical property of $\phi$-divergence is proved in Nguyen2010-xb. In convex optimization (rockafellar1997convex), the subdifferential of a convex function $\phi$ at point $t\in\mathbb{R}$ is defined as the set
@@
We reiterate Lemma 1 in Nguyen2010-xb since it is the foundation of our estimation procedure.
Lemma 1 (Nguyen2010-xb).
For any $\phi$-divergence, the dual form can be written as an optimization problem with respect to a function class $\mathcal{F}$ mapping from $\mathcal{X}$ to $\mathbb{R}$,
@@
This result indicates that the optimizer of the variational form of $\phi$ divergences is a certain function of the density ratio $g$, leading to many ways of estimating the density ratio. Table 1 shows the optimizers for KL-divergence, chi-squared divergence, and the squared Hellinger divergence. In fact, this idea has been explored by nowozin2016f to use the density ratio as the discriminator in a GAN (Generative Adversarial Network).
@@ 2.2 A neural network model to estimate the (relative) density ratio
When $X_{1},\dots,X_{n}\overset{\scriptsize\mbox{iid}}{\sim}P$, and $Y_{1},\dots,Y_{m}\overset{\scriptsize\mbox{iid}}{\sim}Q$, for any measurable set $A\subset\mathcal{X}$, denote the empirical distributions $P_{n}(A):=\frac{1}{n}\sum_{i=1}^{n}\bm{1}\left\{X_{i}\in A\right\}$, and $Q_{m}(A):=\frac{1}{m}\sum_{j=1}^{m}\bm{1}\left\{Y_{j}\in A\right\}$. Using a change-of-variable transformation $g=(1-2f)^{-2}$ from the variational form of the squared Hellinger distance in Table 1, let $\hat{g}_{n}$ be the estimated maximizer over a function class of candidates $\mathcal{G}$, based on the observed samples,
@@
The estimation method depends on the class of candidate functions in $\mathcal{G}$ and the smoothness of the true function $g_{0}$. awaya2025two considers additive tree ensembles for $\log g$ as their $\mathcal{G}$. Note that $g^{*}=p/q$ is the unique global minimizer of $\frac{1}{2}\int g^{-\frac{1}{2}}\text{d}P+\frac{1}{2}\int g^{\frac{1}{2}}\text{d}Q$. In this work, we focus on using neural networks to estimate the ratio functions under this loss through gradient-based optimization. In practice, (stochastic) gradient descent algorithms to find the optimal $\hat{g}_{n},\hat{r}_{n}$.
The feedforward neural network can be viewed as a function learner (Sazli2006BriefReview), parameterized with weight matrices $\left\{W^{(l)}\right\}_{l=1}^{L}$ and bias vectors $\left\{b^{(l)}\right\}_{l=1}^{L}$. Let $U_{l}$ be the total number of nodes in layer $l$ when $1<l\leq L$, $W^{(l)}\in\mathbb{R}^{U_{l}\times U_{l-1}}$, and the vector-valued activation function $\sigma^{(l)}:\mathbb{R}^{U_{l}}\to\mathbb{R}^{U_{l}}$ is an element-wise Lipschitz function. Denote $\bm{\theta}:=\left\{W^{(l)},b^{(l)}\right\}_{l=1}^{L}$ as the set of all parameters, a neural network estimator can be formulated as a sequence of compositions
@@
The empirical loss being minimized was first introduced in awaya2025two and referred to as the “balancing loss” for reasons explained in that paper. The optimizer in (3) can be estimated by $\hat{g}_{n}(x)=f(x;\widehat{\bm{\theta}}_{g})$. Similarly, given the objective in (4) and the learned $\widehat{\bm{\theta}}_{r}$, $\hat{r}_{n}=f(x;\widehat{\bm{\theta}}_{r})$. The neural network architecture varies in different data applications: for image data, variations of convolutional neural networks (CNNs) are applied, and for lower-dimensional examples and microbiome data, multilayer perceptrons (MLPs) are sufficient. Later in Section 3, we will first provide a general convergence theory for general estimators with constraints on $\mathcal{G}$, followed by a convergence rate result using the sieved sparse neural networks space as the function class $\mathcal{G}_{n}$, where the size of the neural network space is allowed to grow with the sample size $n$.